• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Bertin

    Log entry time 15:56:14 on January27,2002

    Entry number 81024

    keyword=Shift summary

    Started shift in kinematics 4G.
    LH2 tasrget with radiator.

    5:00 to 9:00 beam is off, because of the vacuum problem at accelerator ring.
    8:00 had to reboot Left arm magnets slow control IOC.
    8:30 gas alarm, called Jack.
    9:00 dipole fixed back, the hall in beam permit.
    It is imposible to read the point mark on the flore we asume 18.5
    9:25 started data taking at 4GeV.
    11:45 asked controlled access for kinematic 4E
    12:48 Callorimetrer detector is moved for 4E kinematics.
    13:33 beam back, started optics run.
    14 change from otic run to production run. problemes with beam quality,
    position,feedback.
    ask for 20 micoA for the production run
    during run 1780 the target temperature encresed slowly. we keep taking data
    .15h40 go in the kinematic 4Ee change hsr.